# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v2

@@ IV.4.2 Quasiperiodic Model
[…]
We show results for $\|\hat{M}^{(1)}\|_{F}$ in the interacting quasiperiodic model in Fig. 15. For high potential strengths and large system sizes, the conclusions here are largely the same as those in the disordered case. In panels (c) and (d), we see that the distributions of $\|M^{(1)}\|_{F}$ are converged at large $L$, indicating that the transport remains suppressed in the quasiperiodic model for high disorder.
@@ V Conclusion
Many open questions remain. Firstly, our analysis is limited to first-order $(l=1)$ in perturbation theory– it would be interesting to study what happens at higher orders $l$. A challenge arises that with higher-order terms, the resonances become increasingly complex. The second-order term would involve six energy resonances, the third would have eight energy resonances, and so on. Is it possible that this series of corrections converges towards the MBL LIOMs for the single-particle LIOMs with small corrections and large enough disorder? This could be an interesting tool for constructing LIOMs or studying them analytically. For high enough disorder, we conjecture that this perturbation theory will extend to all orders and imply stability of MBL. For low disorder, these increasingly complex resonances may result in large higher-order corrections for increasingly larger fractions of the original noninteracting LIOMs. This would signify a breakdown in perturbation theory for low disorder if one wants to look past the first-order quasi-LIOMs, potentially supporting avalanche instability of the MBL in this regime while still exhibiting parametrically slow relaxation as discussed in Sec. III.4.3 and Sec. IV.5. On a similar note, we also leave open the question of the fate of LIOMs whose corrections are large and suffer from resonances. Here, we do not provide any special treatment for resonating regions of the Hamiltonian, such as what was done in Ref. [76]. It is possible that something similar can be done here so that the perturbation theory can be performed on nonresonating regions only, ensuring its success.
@@
Another open question our analysis can help address is the existence of MBL in $d>1$ dimensions. The seminal work of Ref. [61] argued that avalanches will inevitably destroy MBL in 2D and above. Despite this, experiments [16, 81] and numerics [82, 83, 84] have uncovered signatures of localization in 2D similar to those in the 1D case. Furthermore, the avalanche arguments reconcile poorly with the perturbative arguments presented in Ref. [31], which predicted the existence of MBL in any spatial dimension. It is fair to say that the existence of MBL in $d>1$ is still contested. However, MBL in 2D is extremely difficult to study on all fronts. Signatures obtained in experiments can be difficult to distinguish from glassy phenomena, which can have similar experimental signatures [85]. Furthermore, numerical studies have been largely impossible due to unfavorable scaling with system sizes. The analytical framework presented in this work can, in principle, be applied to any number of dimensions. In particular, one may ask if this approach can be used to study quasiconserved quantities in interacting models in 2D and above with nearest-neighbor interactions. The advantage of our method is the fact that larger system sizes are attainable (relative to what is achievable with exact diagonalization), and calculations scale as a polynomial with the number of sites $L$, as $O(L^{4})$. In this work, we calculated quantities up to 100 sites (which could correspond to a $\times 10$ 10 lattice in 2D), but it is possible to go to larger sizes if one has the computational resources and time. It would be extremely interesting to compare the behavior of quasiconserved quantities in the 2D case to those in the 1D case, as this may shed light on open questions about MBL in 2D. The generalization of the charge transport capacity operator $\hat{M}$ to higher dimensions is nontrivial, but one could study a similar object where the current is defined through a general $(d-1)$-dimensional surface cutting the $d$-dimensional lattice into two partitions. In 1D, this 0-dimensional surface is simply a single link, which was studied in detail in this work. In 2D, one would look at the current through a 1D line. Such a higher-dimensional charge transport capacity object could be an interesting probe for localization in higher dimensions. If one finds that the first-order corrections to LIOMs in 2D are not finite or the charge–transport–capacity–type arguments for bottlenecks do not work, that could strongly hint that localization indeed breaks down in 2D.
@@ Appendix C Additional Results for the AGP
Figure 16: Calculated probability density functions of the logarithm of the Frobenius norm of the AGP, $\ln\|\hat{X}\|_{F}$, similar to Fig. 1 (repeating panel with $W=2.1$ for easy comparisons here), for the weakly-interacting Anderson model. Each distribution was generated absed on 500 disorder realizations with open boundary conditions. Each panel represents a different disorder strength $W$, where disorders from $W=0.6$ to $W=3.1$ are shown, in steps of 0.50.50.5. Within each panel, each curve represents a distribution for a different system size. Here, we show $L=20$ to $L=100$ in steps of 101010.
@@ D.1 Uniform Distribution of Eigenvalues
[…]
@@
Figure 18: Probability density function of $x\equiv 1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\gamma}-\epsilon_{\delta})$ calculated from Eq. (56). For this plot, $\epsilon_{\alpha}$ are assumed to have uniform distributions between 1-1- 1 and 111.
@@ D.2 Density of States of the Quasiperiodic Model
In our work, we observe slower tails ($x^{-1.5}$ and $x^{-1.75}$) in various quantities associated with the quasiperiodic model, when compared to the Anderson model. Similar to how the $\frac{1}{x^{2}}$ tails in the Anderson model can be traced back to the distribution of energy denominators, a comparable mechanism is at play here. However, in the quasiperiodic case, the distribution of combinations like $\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\gamma}-\epsilon_{\delta}$ is highly nontrivial, making it difficult to carry out an analysis analogous to that in Appendix D. Therefore, for reference, we present the density of states (DOS) for the single-particle quasiperiodic model in Fig. 19, computed at $k=\sqrt{2}$ and $L=100$, for values of $h$ below, at, and above the transition point $h=2$. These DOS plots are generated over the spectra of 500 uniformly spaced phase offsets $\delta$ in the interval $[0,\frac{1}{k}]$. The energy levels for a given $\delta$ follow a devil’s staircase-like structure (Cantor function) [87], resulting in a highly complex and nontrivial density of states. This complex structure in the DOS likely underlies the emergence of the observed form of the slower-decaying tails.
Figure 19: Calculated density of states of the quasiperiodic model, where onsite potential values are given by Eq. (4). Here, we fix $k=\sqrt{2}$ and $L=100$. The DOS is calculated over the whole spectrum of each of $\hat{H}$ with 500500500 different phases $\delta$ between $\delta\in\left[0,\frac{1}{k}\right]$.
@@ Appendix E Participation Ratio Data
Figure 20: (a) Median participation ratio (PR) plotted against $L$. The median is taken over 500 disorder realizations, where $\mathrm{PR}$ is calculated for the eigenstate in the middle of the spectrum for each of the disorder realizations. Each colored curve represents a different disorder strength $W$. The dashed line represents the best fit of the form $\mathrm{PR}(L)=A-\frac{A}{1-BL^{\alpha}}$, where $A$ represents the converged $\mathrm{PR}$ value, $\lim_{L\rightarrow\infty}\mathrm{med(PR)}$, and $B,\alpha$ are fitting parameters. (b) Fitted values of $A(W)=\lim_{L\rightarrow\infty}\mathrm{med(PR)}$ versus the disorder strength $W$. A naive polynomial fit to this curve gives a power-law description $\text{PR}\sim W^{-1.78}$.
@@
$\text{PR}=\frac{1}{\sum_{j}|\phi_{\alpha}(j)|^{4}}\leavevmode\nobreak\ .$ (59)
For an orbital with a large localization length in 1D, $\text{PR}\sim\xi$ 999 The participation ratio, while a proxy for the localization length, is not equivalent to it. The localization length $\xi$ is used to characterize the exponential decay of the orbitals $\phi_{i}$, where $\phi_{i}(j)\sim\exp[-|i-j|/\xi]$ On the other hand, the participation ratio measures the extent of the wavefunction $\phi$. Extracting $\xi$ directly from the PR is not straightforward—see, for instance, Ref. [88]. For our purposes, it is sufficient to look at the participation ratio rather than trying to directly calculate $\xi$..
For quick reference, in this section, we show data on how the median of the participation ratio grows with system size in Fig. 20(a), at different disorder strengths $W$. Here, the median is taken over 500500500 disorder realizations, with the participation ratio being calculated for the eigenstate in the middle of the spectrum for each disorder realization. We are particularly interested in the behavior of the $\mathrm{PR}$ in the thermodynamic limit $L\rightarrow\infty$. At small $L$, the curves exhibit an initial power-law growth before saturating to a disorder-dependent plateau that decreases with increasing $W$. Moreover, the PR must vanish as $L\rightarrow 0$. These features motivate the use of the following fitting form for each $\mathrm{med[PR]}$
@@ Appendix G General Bounds on the Charge Transport Capacity
for an arbitrary constant $s$, and $\hat{N}_{\text{tot}}=\sum_{j}\hat{n}_{j}$. This follows from Eq. (27) using that $[\hat{N}_{\text{tot}}]_{\text{off-diag}}=0$, because of the total charge conservation $[\hat{N}_{\text{tot}},\hat{H}]=0$. The Frobenius norm $\|\hat{N}_{R}-s\hat{N}_{\text{tot}}\|_{F}$ can be computed as described in Appendix B for arbitrary bilinear operators:
@@
However, this bound on the operator norm is not very useful in practice, as it is looser than (or equivalent to, for $i_{0}=L/2)$ the more intuitive bound101010 This bound can be obtained by noting that $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\frac{L-i_{0}}{2}\hat{I}\|_{\text{op}}=L-i_{0},$ (77) and $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\hat{N}+\frac{i_{0}}{2}\hat{I}\|_{\text{op}}=i_{0},$ (78) where $\hat{I}$ is the identity operator.
@@ Appendix H Details About the Charge Transport Capacity Matrix Elements and Norms in the Noninteracting Case
$\epsilon_{\alpha}\phi_{j\alpha}=h_{j}\phi_{j\alpha}-\phi_{j+1,\alpha}-\phi_{j-1,\alpha}\leavevmode\nobreak\ ,\quad j=1,\dots,L\leavevmode\nobreak\ .$ (81)
@@
$-\left[\phi_{j-1,\alpha}\phi_{j\beta}-\phi_{j-1,\beta}\phi_{j\alpha}\right]\leavevmode\nobreak\ .$ (82)
@@
$\left(M_{j-1}^{(0)}\right)_{\alpha\beta}-\left(M_{j}^{(0)}\right)_{\alpha\beta}=\phi_{j\alpha}\phi_{j\beta}\leavevmode\nobreak\ .$ (83)
@@
$\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}=-\sum_{j=1}^{i_{0}}\phi_{j\alpha}\phi_{j\beta}\leavevmode\nobreak\ .$ (84)
@@
$\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}=\sum_{j=i_{0}+1}^{L}\phi_{j\alpha}\phi_{j\beta}\leavevmode\nobreak\ .$ (85)
@@
$\qquad\Bigg{(}\sum_{j=i_{0}+1}^{L}\phi_{j\alpha}^{2}\Bigg{)}^{1/2}\Bigg{(}\sum_{j=i_{0}+1}^{L}\phi_{j\beta}^{2}\Bigg{)}^{1/2}\Bigg{\}}\leavevmode\nobreak\ .$

### v3

@@ IV.4.2 Quasiperiodic Model
Figure 16: Numerical computations of the Frobenius norm of the first-order corrections to the charge transport capacity, $\|\hat{M}^{(1)}\|_{F}$, for the weakly-interacting quasiperiodic model with $i_{0}=\lfloor L/2\rfloor$ and open boundary conditions. Each calculation was done based on 500 phase shifts $\delta$ uniformly spaced between 0 and $1/k$. (a)–(d): Calculated probability density functions of $\ln(\|\hat{M}^{(1)}\|_{F})$. Each panel shows a different potential depth $h=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size, between $L=20$ and $L=100$, in steps of 10. The inset in (d) shows a distribution of $\|\hat{M}^{(1)}\|_{F}$ at $h=3.1$ with a log-scale on the y and x-axes, with a $\|\hat{M}^{(1)}\|_{F}^{-1.5}$ (red line) comparison to the tail— see the text surrounding Eq. (24) for details on the form plotted. (e): Plot of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as a function of $L$. Each curve represents a different potential depth $h$, from $h=0.6$ to $3.1$ in steps of 0.5. The inset in (e) shows the same plot but with a log-log scale.
We show results for $\|\hat{M}^{(1)}\|_{F}$ in the interacting quasiperiodic model in Fig. 16. Again, we omit results for $h=0.1$ for visual clarity in the plots. For high potential strengths and large system sizes, the conclusions here are largely the same as those in the disordered case. In panels (c) and (d), we see that the distributions of $\|M^{(1)}\|_{F}$ are converged at large $L$, indicating that the transport remains suppressed in the quasiperiodic model for high disorder.
@@ V Conclusion
Many open questions remain. Firstly, our analysis is limited to first-order $(l=1)$ in perturbation theory– it would be interesting to study what happens at higher orders $l$. A challenge arises that with higher-order terms, the resonances become increasingly complex. The second-order term would involve six energy resonances, the third would have eight energy resonances, and so on. Is it possible that this series of corrections converges towards the MBL LIOMs for the single-particle LIOMs with small corrections and large enough disorder? This could be an interesting tool for constructing LIOMs or studying them analytically. For high enough disorder, we conjecture that this perturbation theory will extend to all orders and imply stability of MBL. For low disorder, these increasingly complex resonances may result in large higher-order corrections for increasingly larger fractions of the original noninteracting LIOMs. This would signify a breakdown in perturbation theory for low disorder if one wants to look past the first-order correction to the LIOMs, potentially supporting avalanche instability of the MBL in this regime while still exhibiting parametrically slow relaxation as discussed in Sec. III.4.4 and Sec. IV.5. On a similar note, we also leave open the question of the fate of LIOMs whose corrections are large and suffer from resonances. Here, we do not provide any special treatment for resonating regions of the Hamiltonian, such as what was done in Ref. [78]. It is possible that something similar can be done here so that the perturbation theory can be performed on nonresonating regions only, ensuring its success.
@@
Another open question our analysis can help address is the existence of MBL in $d>1$ dimensions. The seminal work of Ref. [61] argued that avalanches will inevitably destroy MBL in 2D and above. Despite this, experiments [16, 86] and numerics [87, 88, 89] have uncovered signatures of localization in 2D similar to those in the 1D case. Furthermore, the avalanche arguments reconcile poorly with the perturbative arguments presented in Ref. [31], which predicted the existence of MBL in any spatial dimension. It is fair to say that the existence of MBL in $d>1$ is still contested. However, MBL in 2D is extremely difficult to study on all fronts. Signatures obtained in experiments can be difficult to distinguish from glassy phenomena, which can have similar experimental signatures [90]. Furthermore, numerical studies have been largely impossible due to unfavorable scaling with system sizes. The analytical framework presented in this work can, in principle, be applied to any number of dimensions. In particular, one may ask if this approach can be used to study quasiconserved quantities in interacting models in 2D and above with nearest-neighbor interactions. The advantage of our method is the fact that larger system sizes are attainable (relative to what is achievable with exact diagonalization), and calculations scale as a polynomial with the number of sites $L$, as $O(L^{4})$. In this work, we calculated quantities up to 100 sites (which could correspond to a $10\times 10$ lattice in 2D), but it is possible to go to larger sizes if one has the computational resources and time. It would be extremely interesting to compare the behavior of quasiconserved quantities in the 2D case to those in the 1D case, as this may shed light on open questions about MBL in 2D. The generalization of the charge transport capacity operator $\hat{M}$ to higher dimensions is nontrivial, but one could study a similar object where the current is defined through a general (d1)(d-1)-dimensional surface cutting the $d$-dimensional lattice into two partitions. In 1D, this 0-dimensional surface is simply a single link, which was studied in detail in this work. In 2D, one would look at the current through a 1D line. Such a higher-dimensional charge transport capacity object could be an interesting probe for localization in higher dimensions. If one finds that the first-order corrections to LIOMs in 2D are not finite or the charge–transport–capacity–type arguments for bottlenecks do not work, that could strongly hint that localization indeed breaks down in 2D.
@@ Appendix C Additional Results for the AGP
Figure 17: Calculated probability density functions of the logarithm of the Frobenius norm of the AGP, $\ln\|\hat{X}\|_{F}$, similar to Fig. 1 (repeating panel with $W=2.1$ for easy comparisons here), for the weakly-interacting Anderson model. Each distribution was generated based on 500 disorder realizations with open boundary conditions. Each panel represents a different disorder strength $W$, where disorders from $W=0.6$ to $W=3.1$ are shown, in steps of $0.5$. Within each panel, each curve represents a distribution for a different system size. Here, we show $L=20$ to $L=100$ in steps of $10$.
@@ D.1 Uniform Distribution of Eigenvalues
$P_{\epsilon_{\alpha}+\epsilon_{\beta}}(z)=\int_{x\in X}P_{\epsilon_{\alpha}}(x)P_{\epsilon_{\beta}}(z-x)dx,$ (58)
@@
Figure 19: Probability density function of $x\equiv 1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\gamma}-\epsilon_{\delta})$ calculated from Eq. (59). For this plot, $\epsilon_{\alpha}$ are assumed to have uniform distributions between 1-1 and $1$.
@@ D.2 Density of States of the Quasiperiodic Model
In our work, we observe slower tails ($x^{-1.5}$ and $x^{-1.75}$) in various quantities associated with the quasiperiodic model, when compared to the Anderson model. Similar to how the $\frac{1}{x^{2}}$ tails in the Anderson model can be traced back to the distribution of energy denominators, a comparable mechanism is at play here. However, in the quasiperiodic case, the distribution of combinations like $\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\gamma}-\epsilon_{\delta}$ is highly nontrivial, making it difficult to carry out an analysis analogous to that in Appendix D. Therefore, for reference, we present the density of states (DOS) for the single-particle quasiperiodic model in Fig. 20, computed at $k=\sqrt{2}$ and $L=100$, for values of $h$ below, at, and above the transition point $h=2$. These DOS plots are generated over the spectra of 500 uniformly spaced phase offsets $\delta$ in the interval $0,\frac{1}{k}]$. The energy levels for a given $\delta$ follow a devil’s staircase-like structure (Cantor function) [92], resulting in a highly complex and nontrivial density of states. This complex structure in the DOS likely underlies the emergence of the observed form of the slower-decaying tails.
Figure 20: Calculated density of states of the quasiperiodic model, where onsite potential values are given by Eq. (4). Here, we fix $k=\sqrt{2}$ and $L=100$. The DOS is calculated over the whole spectrum of each of $\hat{H}$ with $500$ different phases $\delta$ between $\delta\in\left[0,\frac{1}{k}\right]$.
@@ Appendix E Participation Ratio Data
Figure 21: (a) Median participation ratio (PR) plotted against $L$. The median is taken over 500 disorder realizations, where $\mathrm{PR}$ is calculated for the eigenstate in the middle of the spectrum for each of the disorder realizations. Each colored curve represents a different disorder strength $W$. The dashed line represents the best fit of the form $\mathrm{PR}(L)=A-\frac{A}{1-BL^{\alpha}}$, where AA represents the converged $\mathrm{PR}$ value, $\lim_{L\rightarrow\infty}\mathrm{med(PR)}$, and $B,\alpha$ are fitting parameters. (b) Fitted values of $A(W)=\lim_{L\rightarrow\infty}\mathrm{med(PR)}$ versus the disorder strength $W$. A naive polynomial fit to this curve gives a power-law description $\text{PR}\sim W^{-1.78}$.
@@
$\text{PR}=\frac{1}{\sum_{j}|\phi_{\alpha}(j)|^{4}}\penalty 10000\ .$ (62)
For an orbital with a large localization length in 1D, $\text{PR}\sim\xi$  [^10] The participation ratio, while a proxy for the localization length, is not equivalent to it. The localization length $\xi$ is used to characterize the exponential decay of the orbitals $\phi_{i}$, where $\phi_{i}(j)\sim\exp[-|i-j|/\xi]$ On the other hand, the participation ratio measures the extent of the wavefunction $\phi$. Extracting $\xi$ directly from the PR is not straightforward—see, for instance, Ref. [93]. For our purposes, it is sufficient to look at the participation ratio rather than trying to directly calculate $\xi$..
For quick reference, in this section, we show data on how the median of the participation ratio grows with system size in Fig. 21(a), at different disorder strengths $W$. Here, the median is taken over $500$ disorder realizations, with the participation ratio being calculated for the eigenstate in the middle of the spectrum for each disorder realization. We are particularly interested in the behavior of the $\mathrm{PR}$ in the thermodynamic limit $L\rightarrow\infty$. At small $L$, the curves exhibit an initial power-law growth before saturating to a disorder-dependent plateau that decreases with increasing $W$. Moreover, the PR must vanish as $L\rightarrow 0$. These features motivate the use of the following fitting form for each $\mathrm{med[PR]}$
@@ Appendix G General Bounds on the Charge Transport Capacity
for an arbitrary constant $s$, and $\hat{N}_{\text{tot}}=\sum_{j}\hat{n}_{j}$. This follows from Eq. (30) using that $\hat{N}_{\text{tot}}]_{\text{off-diag}}=0$, because of the total charge conservation $\hat{N}_{\text{tot}},\hat{H}]=0$. The Frobenius norm $\|\hat{N}_{R}-s\hat{N}_{\text{tot}}\|_{F}$ can be computed as described in Appendix B for arbitrary bilinear operators:
@@
However, this bound on the operator norm is not very useful in practice, as it is looser than (or equivalent to, for $i_{0}=L/2)$ the more intuitive bound [^11] This bound can be obtained by noting that $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\frac{L-i_{0}}{2}\hat{I}\|_{\text{op}}=L-i_{0},$ (80) and $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\hat{N}+\frac{i_{0}}{2}\hat{I}\|_{\text{op}}=i_{0},$ (81) where $\hat{I}$ is the identity operator.
@@ Appendix H Details About the Charge Transport Capacity Matrix Elements and Norms in the Noninteracting Case
$\epsilon_{\alpha}\phi_{j\alpha}=h_{j}\phi_{j\alpha}-\phi_{j+1,\alpha}-\phi_{j-1,\alpha}\penalty 10000\ ,\quad j=1,\dots,L\penalty 10000\ .$ (84)
@@
$-\left[\phi_{j-1,\alpha}\phi_{j\beta}-\phi_{j-1,\beta}\phi_{j\alpha}\right]\penalty 10000\ .$ (85)
@@
$\left(M_{j-1}^{(0)}\right)_{\alpha\beta}-\left(M_{j}^{(0)}\right)_{\alpha\beta}=\phi_{j\alpha}\phi_{j\beta}\penalty 10000\ .$ (86)
@@
$\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}=-\sum_{j=1}^{i_{0}}\phi_{j\alpha}\phi_{j\beta}\penalty 10000\ .$ (87)
@@
$\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}=\sum_{j=i_{0}+1}^{L}\phi_{j\alpha}\phi_{j\beta}\penalty 10000\ .$ (88)
@@
$\qquad\Bigg(\sum_{j=i_{0}+1}^{L}\phi_{j\alpha}^{2}\Bigg)^{1/2}\Bigg(\sum_{j=i_{0}+1}^{L}\phi_{j\beta}^{2}\Bigg)^{1/2}\Bigg\}\penalty 10000\ .$